# Learning large softmax mixtures with warm start EM

### v1

@@ Appendix C Concentration inequalities for quantities related with the random embedding matrix
The following subsections contain deviation inequalities between $N_{\bm{\theta}}$, $\mathrm{I}_{\bm{\theta}}$, $\mathrm{II}_{\bm{\theta}}$, $H_{\bm{\theta}}$ in Eqs. 52 and 53 and their corresponding expectations.
C.1 Concentration inequalities related with $\mathrm{I}_{\theta}$, $\mathrm{II}_{\theta}$ and $N_{\theta}$
@@
Let ${\bm{\theta}}\in\mathbb{R}^{L}$ with $\|{\bm{\theta}}\|_{2}\leq B$ and $s\geq 2$ be arbitrary and assume $p$ is large enough such that
@@
In particular, for arbitrary $\delta>0$, we have for $p$ large enough such that
$\delta^{2}\log p\geq 2s\overline{\sigma}^{2}B^{2}$
@@
Here, $\mathcal{N}_{L}(1/3)$ is a $(1/3)$-net of $\mathbb{S}^{L-1}$ and satisfies $|\mathcal{N}_{L}(1/3)|\leq 7^{L}$ (see, for instance, Vershynin (2018)). Next, we use a truncation device. For fixed ${\bm{\theta}}\in\mathbb{R}^{L}$ with $\|{\bm{\theta}}\|_{2}\leq B$, the random variables $X_{j}^{\top}{\bm{\theta}}$ are zero mean Gaussian random variables with variances $\textrm{Var}({\bm{\theta}}^{\top}X_{j})\leq\|{\bm{\theta}}\|_{2}^{2}\,\overline{\sigma}^{2}\leq B^{2}\overline{\sigma}^{2}$. Consequently, the events
@@
Since $X_{j}^{\top}{\bm{\theta}}$ is Gaussian, the distribution of $(X_{j}^{\top}{\bm{\theta}})^{2}$ is sub-exponential (with parameter $\leq\overline{\sigma}^{2}$). This implies that the distribution of
@@
Let ${\bm{\theta}}\in\mathbb{R}^{L}$ with $\|{\bm{\theta}}\|_{2}\leq B$ and $\delta>0$ and $s\geq 2$. For $p\geq p_{0}=p_{0}(B,\delta,s,\overline{\sigma})$, we have
$\mathbb{P}\left\{\left\|\mathrm{II}_{\bm{\theta}}-\mathbb{E}[\mathrm{II}_{\bm{\theta}}]\right\|_{2}\lesssim\overline{\sigma}\sqrt{L+\log(p)}p^{\frac{1}{2}+\delta}\right\}\geq 1-4p^{1-s}.$
@@
Let ${\bm{\theta}}\in\mathbb{R}^{L}$ with $\|{\bm{\theta}}\|_{2}\leq B$ and $\delta>0$ and $s\geq 2$. For $p\geq p_{0}=p_{0}(B,\delta,s,\overline{\sigma})$, we have
@@
Again, we follow the same arguments that we used to prove Lemmas 8 and 9. Recall the events $\{\mathcal{X}_{j}(s,{\bm{\theta}})\}_{j\in[p]}$ from (122). On the intersection of the events, we have
$\left|N_{\bm{\theta}}-\mathbb{E}[N_{\bm{\theta}}]\right|$ $\leq\left|\sum_{j=1}^{p}\left(e^{X_{j}^{\top}{\bm{\theta}}}1_{\mathcal{X}_{j}(s,{\bm{\theta}})}-\mathbb{E}\left[e^{X_{j}^{\top}{\bm{\theta}}}1_{\mathcal{X}_{j}(s,{\bm{\theta}})}\right]\right)\right|+\sum_{j=1}^{p}\mathbb{E}\left[e^{X_{j}^{\top}{\bm{\theta}}}1_{\mathcal{X}_{j}^{c}(s,{\bm{\theta}})}\right]$
@@
C.2 Deviation inequalities of $\mathrm{I}_{\theta}$, $\mathrm{II}_{\theta}$ and $N_{\theta}$
The following lemma contains results on the moments related with $\mathrm{I}_{\bm{\theta}}$, $\mathrm{II}_{\bm{\theta}}$ and $N_{\bm{\theta}}$.
@@ C.2 Deviation inequalities of $\mathrm{I}_{\theta}$, $\mathrm{II}_{\theta}$ and $N_{\theta}$
Let ${\bm{\theta}}\in\mathbb{R}^{L}$. Let $N_{\bm{\theta}}$, $\mathrm{I}_{\bm{\theta}}$ and $\mathrm{II}_{\bm{\theta}}$ be defined in (52). For any $v\in\mathbb{R}^{L-1}$, we have
@@
The following lemma follows immediately from Lemmas 8, 9 & 10.
@@
Let ${\bm{\theta}}\in\mathbb{R}^{L}$ with $\|{\bm{\theta}}\|_{2}\leq B$, $s\geq 2$, $\delta>0$ and $\epsilon>0$. For $p\geq p_{0}(B,s\overline{\sigma},\delta,\epsilon)$, the following holds with probability at least $1-4p^{1-s}$:
@@
C.3 Concentration inequalities related with $H_{\theta}$
For any ${\bm{\theta}}\in\mathbb{R}^{L}$, define
$\bar{H}_{\bm{\theta}}={\mathbb{E}[\mathrm{I}_{\bm{\theta}}]\over\mathbb{E}[N_{\bm{\theta}}]}-{\mathbb{E}[\mathrm{II}_{\bm{\theta}}]\mathbb{E}[\mathrm{II}_{\bm{\theta}}]^{\top}\over(\mathbb{E}[N_{\bm{\theta}}])^{2}}.$
The following lemma is a key result that provides deviation inequality of $\|H_{\bm{\theta}}-\bar{H}_{\bm{\theta}}\|_{\mathrm{op}}$.
@@ C.3 Concentration inequalities related with $H_{\theta}$
Let ${\bm{\theta}}\in\mathbb{R}^{L}$ with $\|{\bm{\theta}}\|_{2}\leq B$, $s\geq 2$, $\delta>0$ and $\epsilon>0$. For $p\geq p_{0}(B,s,\overline{\sigma},\delta,\epsilon)$,
$\mathbb{P}\left\{\|H_{\bm{\theta}}-\bar{H}_{\bm{\theta}}\|_{\mathrm{op}}\lesssim\overline{\sigma}^{2}\sqrt{L+\log(p)\over p^{1-\delta}}\right\}\geq 1-4p^{1-s}$ (126)
Moreover, for any $\epsilon>0$, with the same probability as above, we have

### v2

@@ Appendix D Concentration inequalities for quantities related with the random embedding matrix
The following subsections contain deviation inequalities between $N_{\bm{\theta}}$, $\mathrm{I}_{\bm{\theta}}$, $\mathrm{II}_{\bm{\theta}}$, $H_{\bm{\theta}}$ in Eqs. 62 and 63 and their corresponding expectations, derived under the following distributional assumption on the rows of $\mathbf{X}$.
Assumption 7.
The rows of $\mathbf{X}$ are i.i.d. sub-Gaussian random vectors with zero mean and sub-Gaussian constant $\overline{\sigma}<\infty$. [^2] A random vector $Z\in\mathbb{R}^{d}$ is said to be $\gamma$ sub-Gaussian if for any $v\in\mathbb{R}^{d}$, $\mathbb{E}[\exp(v^{\top}Z)]\leq\exp(\|v\|_{2}^{2}\gamma^{2}/2)$.
D.1 Concentration inequalities related with $\mathrm{I}_{\theta}$, $\mathrm{II}_{\theta}$ and $N_{\theta}$
@@
Grant 7. Let ${\bm{\theta}}\in\mathbb{R}^{L}$ with $\|{\bm{\theta}}\|_{2}\leq B$ and $s\geq 2$ be arbitrary and assume $p$ is large enough such that
@@
In particular, for arbitrary $\delta>0$ and for $p$ large enough such that
$\delta^{2}\log p\geq 2s\overline{\sigma}^{2}B^{2},$
we have
@@
Here, $\mathcal{N}_{L}(1/3)$ is a (1/3)(1/3)( 1 / 3 )-net of $\mathbb{S}^{L-1}$ and satisfies $|\mathcal{N}_{L}(1/3)|\leq 7^{L}$ (see, for instance, Vershynin (2018)). Next, we use a truncation device. For fixed ${\bm{\theta}}\in\mathbb{R}^{L}$ with $\|{\bm{\theta}}\|_{2}\leq B$, the random variables $X_{j}^{\top}{\bm{\theta}}$ are zero mean sub-Gaussian random variables with sub-Gaussian constant no greater than $\|{\bm{\theta}}\|_{2}^{2}\,\overline{\sigma}^{2}\leq B^{2}\overline{\sigma}^{2}$. Consequently, the events
@@
Since $X_{j}^{\top}v$ is $\overline{\sigma}$ sub-Gaussian, the distribution of $(X_{j}^{\top}v)^{2}$ is sub-exponential (with parameter $\leq\overline{\sigma}^{2}$). This implies that the distribution of
@@
Grant 7. Let ${\bm{\theta}}\in\mathbb{R}^{L}$ with $\|{\bm{\theta}}\|_{2}\leq B$ and $\delta>0$ and $s\geq 2$. For $p\geq p_{0}=p_{0}(B,\delta,s,\overline{\sigma})$, we have
$\mathbb{P}\left\{\left\|\mathrm{II}_{\bm{\theta}}-\mathbb{E}[\mathrm{II}_{\bm{\theta}}]\right\|_{2}\lesssim\overline{\sigma}\sqrt{L+\log(p)}~p^{\frac{1}{2}+\delta}\right\}\geq 1-4p^{1-s}.$
@@
Grant 7. Let ${\bm{\theta}}\in\mathbb{R}^{L}$ with $\|{\bm{\theta}}\|_{2}\leq B$ and $\delta>0$ and $s\geq 2$. For $p\geq p_{0}=p_{0}(B,\delta,s,\overline{\sigma})$, we have
@@
Again, we follow the same arguments that we used to prove Lemmas 12 and 13. Recall the events $\{\mathcal{X}_{j}(s,{\bm{\theta}})\}_{j\in[p]}$ from (130). On the intersection of the events, we have
$|\displaystyle\left|N_{\bm{\theta}}-\mathbb{E}[N_{\bm{\theta}}]\right|$$\leq\left|\sum_{j=1}^{p}\left(e^{X_{j}^{\top}{\bm{\theta}}}1_{\mathcal{X}_{j}(s,{\bm{\theta}})}-\mathbb{E}\left[e^{X_{j}^{\top}{\bm{\theta}}}1_{\mathcal{X}_{j}(s,{\bm{\theta}})}\right]\right)\right|+\sum_{j=1}^{p}\mathbb{E}\left[e^{X_{j}^{\top}{\bm{\theta}}}1_{\mathcal{X}_{j}^{c}(s,{\bm{\theta}})}\right]$
@@
D.2 Concentration inequalities related with $H_{\theta}$ under Gaussianity
For any ${\bm{\theta}}\in\mathbb{R}^{L}$, the following lemma contains results on the moments related with $\mathrm{I}_{\bm{\theta}}$, $\mathrm{II}_{\bm{\theta}}$ and $N_{\bm{\theta}}$ under the condition
Assumption 8.
@@ D.2 Concentration inequalities related with $H_{\theta}$ under Gaussianity
Grant 8. Let ${\bm{\theta}}\in\mathbb{R}^{L}$. Let $N_{\bm{\theta}}$, $\mathrm{I}_{\bm{\theta}}$ and $\mathrm{II}_{\bm{\theta}}$ be defined in (62). For any $v\in\mathbb{R}^{L-1}$, we have
@@
The following lemma follows immediately from Lemmas 12, 13 & 14.
@@
Grant 8. Let ${\bm{\theta}}\in\mathbb{R}^{L}$ with $\|{\bm{\theta}}\|_{2}\leq B$, $s\geq 2$, $\delta>0$ and $\epsilon>0$. For $p\geq p_{0}(B,s,\overline{\sigma},\delta,\epsilon)$, the following holds with probability at least $1-4p^{1-s}$:
@@
The following lemma is a key result that provides deviation inequality of $\|H_{\bm{\theta}}-\bar{H}_{\bm{\theta}}\|_{\mathrm{op}}$ where
$\bar{H}_{\bm{\theta}}={\mathbb{E}[\mathrm{I}_{\bm{\theta}}]\over\mathbb{E}[N_{\bm{\theta}}]}-{\mathbb{E}[\mathrm{II}_{\bm{\theta}}]\mathbb{E}[\mathrm{II}_{\bm{\theta}}]^{\top}\over(\mathbb{E}[N_{\bm{\theta}}])^{2}}.$ (134)
@@
Grant 8. Let ${\bm{\theta}}\in\mathbb{R}^{L}$ with $\|{\bm{\theta}}\|_{2}\leq B$ and fix any $s\geq 2$, $\delta>0$ and $\epsilon>0$. For $p\geq p_{0}(B,s,\overline{\sigma},\delta,\epsilon)$,
$\mathbb{P}\left\{\|H_{\bm{\theta}}-\bar{H}_{\bm{\theta}}\|_{\mathrm{op}}\lesssim\overline{\sigma}^{2}\sqrt{L+\log(p)\over p^{1-\delta}}\right\}\geq 1-4p^{1-s}.$ (135)
Moreover, with the same probability as above, we have